A reaction has an equilibrium constant of 7.4×103 at 298 K. At 752 K , the equilibrium constant is 0.43.

Find ΔHorxn for the reaction.
Answer #1

Given:

T1 = 298 K

T2 = 752 K

K1 = 7.4*10^3

K2 = 0.43

use:

ln(K2/K1) = (Ho/R)*(1/T1 - 1/T2)

ln(0.43/7.4*10^3) = ( Ho/8.314)*(1/298 - 1/752)

-9.7532 = (Ho/8.314)*(2.026*10^-3)

Ho = -40025 J/mol

Ho = -40.0254 KJ/mol

Answer: -40.0 KJ/mol